MILLIMETER-WAVE SPECTROSCOPY OF THE vdW BANDS OF He-HCN NEAR THE DISSOCIATION LIMIT.

KENSUKE HARADA and \underline{KEIICHI TANAKA}, Department of Chemistry, Faculty of Science, Kyushu University, Hakozaki, Higashiku, Fukuoka, 812-8581 Japan.

The He-HCN complex is a weakly bound complex with binding energy of about 9 cm^{−1}. We have measured the internal rotation bands (to the j = 1 and 2 states) and intermolecular stretching bands of the He-HCN complex by millimeter-wave absorption spectroscopy combined with a pulsed-jet expansion technique, and reported the potential energy surface (PES) to reproduce the observed transition frequencies.^{a} The PES obtained has a global minimum in the linear configuration with a depth of 30.3 cm^{−1} and has a saddle point at the anti-linear configuration with a depth of 20.3 cm^{−1}. The intermolecular stretching level and the second excited (j = 2) level of the internal rotation state are determined to be located 9.1407 and 9.0529 cm^{−1}, respectively, above the ground state, very close to the calculated dissociation limit (D_{0}) of 9.33 cm^{−1} . In the present study, we have extended the measurement to the transitions to the bound states above the ”dissociation limit” (D_{0}) and observed several such transitions for the first time in the ground state of the molecular complex. The rovibrational levels of He-HCN with e label dissociate to the HCN molecule with J = 0 and the He atom (D_{0}), while those with f label, due to the parity conservation, to the HCN molecule with J = 1 and the He atom which is higher in energy by about 2B_{HCN} ∼ 2.96 cm^{−1} than D_{0}, as discussed in the infrared study of He-HF.^{b} The PES obtained in the present analysis indicates that four f levels in the j = 1 and 2 excited states are bound above the ”dissociation limit” (D_{0}) due to the parity conservation. In addition five levels (both of e and f labels) are predicted to be bound by centrifugal barrier with finite life times but not yet observed.

aK. Harada, K. Tanaka, T. Tanaka, S. Nanbu, and M. Aoyagi, J. Chem. Phys. 117, 7041 (2002).

bC. M. Lovejoy and D. J. Nesbitt, J. Chem. Phys. 93, 5387 (1990).